OpenStudy (anonymous):

Find an equation of the line. Write the equation using function notation. Through (8,-3); parallel to 8x+3y=79

OpenStudy (anonymous):

the required line is parallel to 8x+3y=97 so the slopes of the two lines are the same

OpenStudy (anonymous):

for the given line are u can specify the slope?

OpenStudy (anonymous):

Do we need to use the y-intercept form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=mx+b, (8,3) 8x+3y=79

OpenStudy (anonymous):

you have 8x+3y=79 then 3y=-8x+79 and y=-8/3 x+79/3 then by compare this equation with the general one (y=mx+c) m=-8/3

OpenStudy (anonymous):

now you have m=-8/3 and a point (8,-3)

OpenStudy (anonymous):

use the equation \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

8-3

OpenStudy (anonymous):

I'm having a hard time trying figure out the y-y1=m(x-x1)

OpenStudy (anonymous):

ok y+3=-8/3 (x-8) y+3=-8/3 x+64/3 y=-8/3 x +64/3 -3 y=-8/3 x +55/3